OpenStudy (anonymous):

Find the number c that satisfies the conclusion of the Mean Value Theorem. f(x)=e^(-3x) [0,2] Find the number c that satisfies the conclusion of the Mean Value Theorem. f(x)=e^(-3x) [0,2] @Mathematics

OpenStudy (anonymous):

solve \[\frac{e^{-6}-e^0}{2}=-3e^{-3c}\] for c

OpenStudy (anonymous):

Thanks for the reply! I have to head back to the dorm, will look at it in about 30minutes. Thanks a bunch satellite! REALLY appreciate it!

OpenStudy (anonymous):

ok i will write out the first few steps, you can check back later

OpenStudy (anonymous):

first of all you are using \[\frac{f(b)-f(a)}{b-a}=f'(c)\] with \[f(x)=e^{-3x}, f'(x)=-3e^{-3x}, a = 0, b = 2 \] giving \[\frac{e^{-6}-e^0}{2-0}=\frac{e^{-6}-1}{2}\] \[=\frac{1-e^6}{2e^6}\] so you need to solve \[\frac{1-e^6}{2e^6}=-3e^{-3c}\] or \[\frac{1-e^6}{2e^6}=-\frac{3}{e^{3c}}\] and it is algebra from here on in.

OpenStudy (anonymous):

get \[\frac{2e^6}{1-e^6}=-\frac{e^{3c}}{3}\] \[\frac{6e^6}{e^6-1}=e^{3c}\] \[3c=\ln(6)+\ln(e^6)-\ln(e^6-1)\] \[c=\frac{1}{3}(\ln(6)+6-\ln(e^6-1))\]

OpenStudy (anonymous):

how did you get from (e^-6 -1 )/2 to 1-e^-6/2e^6

OpenStudy (anonymous):

multiplied top and bottom by \[e^6\] to get rid of the compound fraction

OpenStudy (anonymous):

it should be \[\frac{e^{-6}-1}{2}=\frac{1-e^6}{2e^6}\] the negative exponents annoyed me, because it is really a compound fraction. you can probably work with them but i like to see what is really going on to solve for c

OpenStudy (anonymous):

Got it. Thanks bud!

OpenStudy (anonymous):

yw